# Class Agnostic Instance-level Descriptor for Visual Instance Search

### v1

@@ V-A Experiment Setup
V-B Instance Search Performance
In the first experiment, the effectiveness of our method is studied in comparison to several representative methods in instance search. Among these methods, R-MAC [17], CAM-weight [41], BLCF [20] and BLCF-SalGAN [20] are deep global features which collapse all the features of an image into one vector for retrieval. In contrast, methods like DASR [4] and ours extract the instance level feature from the candidate images and are able to localize the instance from the candidate images. In addition, we also treat the unsupervised object detector CutLER [10] as an instance region detector. The same feature as ours is built for the regions detected by CutLER. Moreover, we also report the results when the bisecting clustering in CLAID is replaced by NCut [28], while keeping other configurations unchanged. The instance search performance is shown in Table I.
As seen from the table, global features show poor performance on the instance search task. Object level features such as DASR, CutLER, and CLAID demonstrate considerably better performance. CLAID outperforms the rest of methods on three benchmarks. Particularly, CLAID shows considerably better performance than the rest on INSTRE where the queries are more challenging and there are a lot of distractors in the reference set. Although CutLER is also able to build instance level features, its performance on instance search is considerably lower than CLAID. The relatively poor performance mainly dues to its limited number of partitions over the feature set. Moreover, the cut in a chain in CutLER also causes the loss of potential instance regions. Compared to CLAID (NCut), CLAID performs considerably better. According to our observation, the results from bisecting clustering better reflect the multi-scale semantic structure than the partitions produced by NCut. In addition, we also tested different backbones with CLAID at its feature extraction stage. As shown from the table, our method shows considerably better performance when it is supported by the same backbone as the competing methods (e.g., DASR and CutLER). This largely owes to its high localization accuracy. Figure 4 further showcases the instance search results from CLAID.
In order to study the behavior of our method on multi-instance search task, the performance on 250 queries from INSTRE is particularly evaluated. For these queries, there are more than one objects in presence in a query. As shown in Table II, most of the methods show relatively better performance on this task than that of single instance search (reported in Table I). The search quality from all global features is considerably better. This is mainly because the large coverage of the instance query over the image makes it close to a conventional content-based image search task. The mixture of features from different instances becomes helpful in this context. Our method still shows the best performance as it keeps the features extracted from different scales. The localization accuracy of different methods (measured by mIoU) is reported on Table III. Since the localization is impossible to fulfill for the global features, only the performance from instance level features is reported. The patch-level region masks from our method and CutLER are transformed to bounding boxes when we calculate the mIoU score. As shown in Table III, CLAID shows considerably better localization accuracy over DASR* and CutLER.
V-C The Performance of Feature Detector
Figure 5: Recall at different IoU thresholds for four methods on COCO 20k.
Since the feature detector is the critical component in our pipeline, its effectiveness is further studied from two perspectives, namely the detection speed and the exhaustiveness in detecting latent instance regions from an image. The processing speed is compared to unsupervised object detectors such as Deep Spectral [11], TokenCut [12], and CutLER [10]. All of them are built on self-supervised ViT. For this reason, we only evaluate the time cost of processing one image. The speed efficiency is evaluated on CPU with single-thread.
The average number of cuts that are carried out on one image, the average time of processing one image, and the time cost of one cut are shown in Table IV. In TokenCut, CutLER, and CLAID (NCut), NCut [28] is adopted to carry out the partition. The NCut is computationally intensive as it involves the eigenvalue decomposition on a matrix in big size. The time cost of one cut is much higher than that of the bisecting clustering adopted by CLAID. The processing time for CLAID (NCut) is relatively shorter than the other methods based on NCut. The cut time becomes shorter since the cut is applied on the small subsets on the hierarchy. Owing to the efficiency of bisecting clustering, CLAID is 6$\times$ faster than CLAID (Ncut). To verify the ability of our detector in detecting the latent instance regions in an arbitrary image, we further study the performance of CLAID on COCO 20k segmentation validation dataset. In this evaluation, CLAID is treated as an unsupervised instance detection method. The exhaustiveness is measured by the recall rate of the object instances in different categories that have been annotated in the dataset. The recall is evaluated with respect to mIoU score. Three detectors Deep Spectral [11], TokenCut [12] and CutLER [10] are considered in the comparison. The detection performance is shown in Figure 5. As seen from the figure, CLAID outperforms these detectors by a large margin because the hierarchical decomposition is able to discover more correct object regions.
TABLE IV: The comparison on the processing time costs
Methods ##\##cut TM per image (s) TM per cut (s)
@@ V-C The Performance of Feature Detector
CLAID 29 2.40 0.083
V-D Performance on Image Retrieval
Generally speaking, image retrieval can be viewed as a multi-instance search task. Our method is, therefore, also feasible for conventional image retrieval. In this experiment, we study the effectiveness of CLAID on image retrieval. Following the same processing pipeline, a bunch of CLAID features are extracted from one candidate image. While only one feature is extracted from a query image. The ranking list is produced based on the position where a candidate image first appears in the instance search ranking list.
Performance is evaluated on three benchmarks Paris6k [42], Holidays [43], and Oxford5k [44]. As shown from the Table V, CLAID consistently outperforms a majority of representative methods in the literature. The best performance is achieved by the DUODIS [27]. DUODIS excels by mining latent objects directly from the test datasets and employs a pre-trained model, which is pre-conditioned on images of landmarks and buildings for feature extraction.
TABLE V: Performance (mAP) on image retrieval. digits are from the referred paper. digits are from  [4]
@@ V-E Ablation Studies
To study the impact of each major component in CLAID, an ablation analysis is carried out on INSTRE. Namely, we study the impact of seed selection (Section III-B) and the contribution from the filtering scheme on the “dummy node” (Section III-D). Table VI showcases the mAP results and the total number of features extracted from INSTRE images under each configuration. As shown in the Table, compared to CLAID with random initialization, the detector produces 16% less redundant instance regions. Moreover, compared to the run extracting features for dummy nodes, the standard configuration of CLAID shows slightly higher mAP while using 42% less number of features. This basically indicates that regions that correspond to the dummy nodes make no contribution to the instance level feature representation.
TABLE VI: Ablation study on the seed selection and the filtering strategy on INSTRE
@@
V-F Sensitivity Studies
In our method, there are two key hyperparameters namely the average internal connectivity ($\tau_{1}$) and the level of semantic dominance ($\tau_{2}$). To evaluate the impact of hyperparameters settings on our model’s performance, a series of sensitivity tests have been pulled out. The experiments are carried out on Instance-160. In the experiments, when validating $\tau_{2}$, we fix $\tau_{1}$ to 0.97. Conversely, when validating $\tau_{1}$, we fixed $\tau_{2}$ to 0.2. This setup allows us to isolate the impact of each parameter and analyze their individual impact on the model’s performance. The instance search performance trends on Instance-160 are presented in Figure 6. As shown from the figure, the search performance is more sensitive to the level of semantic dominance ($\tau_{2}$), while only minor performance variation is observed when $\tau_{1}$ varies. Nevertheless, the search performance remains stable as long as $\tau_{2}$ is in the range $[0.2,~{}0.5]$.
@@ VI Conclusions
We have presented a novel instance level feature descriptor, namely CLAID. Given an image is represented by the feature set produced by self-supervised ViT, a novel top-down repeated bisecting procedure is proposed to partition the feature set into compact and spatially connected instance regions. The nodes on the resulting hierarchy correspond to salient instance regions mined from the image. Compared to the existing methods, it provides a more exhaustive description about the latent instances in one image. Most importantly, our framework is class-agnostic, which allows it to remain effective on various object categories. Excellent performance is achieved on single and multiple instance search tasks as well as image search task. It, therefore, becomes possible to unify the visual instance search and the content-based image search under one framework. Moreover, only around 30 features are extracted per image, it allows the search to be scaled up to billion images level on a single server.

### v2

@@ V-C The Performance of Feature Detector
V-D Performance on Image Retrieval
Generally speaking, image retrieval can be viewed as a multi-instance search task. Our method is, therefore, also feasible for conventional image retrieval. In this experiment, we study the effectiveness of CLAID on image retrieval. Following the same processing pipeline, a bunch of CLAID features are extracted from one candidate image. Only one feature is extracted from a query image. The ranking list is produced based on the position where a candidate’s image first appears in the instance search ranking list.
Performance is evaluated on three benchmarks Paris6k [48], Holidays [49], and Oxford5k [50]. As shown from the Table VI, CLAID consistently outperforms a majority of representative methods in the literature. The best performance is achieved by the DUODIS [28]. DUODIS excels by mining latent objects directly from the test datasets and employs a pre-trained model, which is pre-conditioned on images of landmarks and buildings for feature extraction.
TABLE V: The comparison on the processing time costs
Methods #\#cut TM per image (s) TM per cut (s)
@@ V-D Performance on Image Retrieval
CLAID 29 2.4 0.083
TABLE VI: Performance (mAP) on image retrieval. digits are from the referred paper. digits are from  [4]
@@ V-E Ablation Studies
TABLE VII: Ablation study on the seed selection and the filtering strategy on INSTRE
@@
TABLE VIII: Instance Search Performance (mAP) when setting the decomposition depth as the termination threshold
Datasets Depth #Features mAP-50 mAP-100 mAP-all
Instance-335 3 572,796 0.705 0.624 0.651
5 2,536,668 0.727 0.647 0.670
7 10,392,156 0.737 0.657 0.678
CLAID - 1,473,165 0.778 0.710 0.725
INSTRE 3 383,102 0.857 0.768 0.686
5 1,692,166 0.872 0.788 0.713
7 6,932,422 0.881 0.801 0.729
CLAID - 827,426 0.894 0.827 0.775
To study the impact of each major component in CLAID, an ablation analysis is carried out on INSTRE. Namely, we study the impact of seed selection (Section III-B) and the contribution from the filtering scheme on the “dummy node” (Section III-D). Table VII shows the mAP results and the total number of features extracted from INSTRE images under each configuration. As shown in the Table, compared to CLAID with random initialization, the detector produces 16% less redundant instance regions. Moreover, compared to the run extracting features for dummy nodes, the standard configuration of CLAID shows slightly higher mAP while using 42% fewer features. This basically indicates that regions that correspond to the dummy nodes make no contribution to the instance-level feature representation.
In order to study the efficacy of the proposed termination condition (defined in Eqn. 9), a variant run of the method is pulled out. In this study, the termination condition is replaced with the decomposition depth. The maximum decomposition depth is varied from 3 to 7. The filtering with “dummy node” condition is disabled as well. The experiments are carried out on INSTRE and Instance-335. The instance search performance is compared to the standard configuration of CLAID and presented in Table VIII. The performance gap to the standard CLAID can be reduced to 3-4% when the images are decomposed into a growing number of fine-grained regions. Nevertheless, several times more features are produced per image.
V-F Configuration Studies
@@
In our method, there are two key hyperparameters, namely the average internal connectivity ($\tau_{1}$) and the level of semantic dominance ($\tau_{2}$). To evaluate the impact of hyperparameter settings on our model’s performance, the sensitivity tests have been pulled out. The experiments are carried out on Instance-160. In the experiments, when validating $\tau_{2}$, we fix $\tau_{1}$ to 0.97. Conversely, when validating $\tau_{1}$, we fixed $\tau_{2}$ to 0.2. This setup allows us to isolate the impact of each parameter and analyze its individual impact on the model’s performance. The instance search performance trends on Instance-160 are presented in Figure 8. As shown from the figure, the search performance is more sensitive to the level of semantic dominance ($\tau_{2}$), while only minor performance variation is observed when $\tau_{1}$ varies. Nevertheless, the search performance remains stable as long as $\tau_{2}$ is in the range [0.2,0.5][0.2,~0.5].
@@ VI Conclusions
We have presented a novel instance-level feature, namely CLAID. A novel top-down bisecting procedure has been introduced to decompose the feature set, which is produced by a pretrained ViT, recursively into a hierarchy. To support the decomposition, an adaptive termination condition and the “dummy node” filtering strategy have been proposed. Each node on the hierarchy corresponds to a potential instance region in the image. As a consequence, we are able to decompose an image into a collection of instance regions in different granularities and represent them with features of uniform length. The features are class-agnostic because the detector is built on non-classification ViT backbone. These features may cover a dominant image region, an integral of multiple instances, or a single instance. It, therefore, supports the similarity measurement between two images as well as two instances. As shown by the comprehensive experiments, excellent performance is achieved on instance search as well as image retrieval. Additional experiments also show the high compatibility of our method with different pretrained ViT backbones, such as DINOv2 and SigLIP. In particular, the excellent performance on SigLIP does indicate the feasibility of our method to unify the text-to-instance search, instance search, and image retrieval into one search framework.